# Loss-Guided Auxiliary Agents for Overcoming Mode Collapse in GFlowNets

### v1

@@ A.1.2 Valid sequences
For this task, we use GPT2Model from Hugging Face’s transformers library [55] with 3 hidden layers, 8 attention heads, and embeddings of dimension 64. For LGGFN, $\lambda$ was set to $1$ and for SAGFN, $\{\beta_{e}^{\text{main}}=1,\beta_{e}^{\text{aux}}=1,\beta^{\text{aux}}=1,\beta^{\text{main}}=1,\beta_{i}=1\}$.
A.1.3 Bayesian Structure Learning
For this task, we use a directional graph-based neural architecture to compute action logits for graph generation. The model supports decisions for both edge creation between nodes and an optional stop action that signals the end of generation. The network begins by embedding node identifiers into a continuous feature space. These embeddings are processed through multiple layers of direction-aware graph convolutions that distinguish between incoming and outgoing edges. Each layer includes residual connections to preserve information across depths and applies layer normalization to stabilize training. The output of each convolutional layer consists of separate representations for incoming and outgoing directional features. These are refined independently using small feedforward networks and then recombined into a unified node representation. To produce edge-level action scores, the model computes dot products between the feature vectors of source and target nodes. In addition, a pooled summary of all node embeddings can be used to compute a score for the optional stop action. The final output consists of unnormalized logits representing the desirability of all possible edge additions and, if applicable, the decision to terminate the graph construction process.
@@ A.1.3 Bayesian Structure Learning
A.2 Visualization of HyperGrid learned distributions
Figure 8: Learnt distributions on environment with size from 4 to 64.

### v2

@@ Conclusion
For this task, we use GPT2Model from Hugging Face’s transformers library (wolf2019huggingface) with 3 hidden layers, 8 attention heads, and embeddings of dimension 64. For LGGFN, $\lambda$ was set to $1$ and for SAGFN, $\{\beta_{e}^{\text{main}}=1,\beta_{e}^{\text{aux}}=1,\beta^{\text{aux}}=1,\beta^{\text{main}}=1,\beta_{i}=1\}$.
# Nodes 5 6 7 8 9 10 11 12 13
Bayesian Structure Learning:
Markov Equivalence:
Two graphs $G_{1}$ and $G_{2}$ are said to be Markov equivalent if they encode the same set of conditional independencies (Figure 12):
$G_{1}\sim G_{2}\;\Leftrightarrow\;CI(G_{1})=CI(G_{2})$
$\text{where}\quad CI(G)=\{(X,Y,Z)\mid X\perp\!\!\!\perp Y\mid Z\text{ in }G\}$
[…]
Figure 12: Three Markov equivalent DAGs.
For this environment, we first generate a random directed acyclic graph (DAG) (by sampling an Erdős–Rényi graph and assigning edge directions such that the source node has a lower index than the target node). Data is then generated according to a linear-Gaussian Structural Causal Model (SCM):
In this setting, the states of the environment correspond to DAGs, while the actions consist of adding valid edges, i.e., those that preserve acyclicity. The reward function for a given graph is defined based on the graph itself and its parent graphs in the state space:
$\log R(G)=\sum_{j=1}^{d}\text{LocalScore}(X_{j}\mid\text{Pa}_{G}(X_{j})),$
where LocalScore denotes the BGe score (heckerman1995learning), and $\text{Pa}_{G}(X_{j})$ represents the set of parents of node $X_{j}$ in graph $G$. For this task, we use a directional graph-based neural architecture to compute action logits for graph generation. The model supports decisions for both edge creation between nodes and an optional stop action that signals the end of generation. The network begins by embedding node identifiers into a continuous feature space. These embeddings are processed through multiple layers of direction-aware graph convolutions that distinguish between incoming and outgoing edges. Each layer includes residual connections to preserve information across depths and applies layer normalization to stabilize training. The output of each convolutional layer consists of separate representations for incoming and outgoing directional features. These are refined independently using small feedforward networks and then recombined into a unified node representation. To produce edge-level action scores, the model computes dot products between the feature vectors of source and target nodes. In addition, a pooled summary of all node embeddings can be used to compute a score for the optional stop action. The final output consists of unnormalized logits representing the desirability of all possible edge additions and, if applicable, the decision to terminate the graph construction process.
@@
mRNA Sequence Design:
Messenger RNA (mRNA) sequences are composed of codons, triplets of nucleotides, each of which encodes a specific amino acid. Due to the redundancy of the genetic code, most amino acids can be encoded by multiple synonymous codons. Thus, a single protein sequence can correspond to many possible mRNA sequences. Leveraging this redundancy, we introduce the CodonDesign environment to model codon-level mRNA sequence design conditioned on a fixed protein sequence. States represent partially constructed mRNA sequences, and actions correspond to choosing synonymous codons or a special exit action. At each position $t$, only codons that encode the $t$-th amino acid are permitted. The episode terminates after $T$ steps (the protein length), followed by the exit action. The reward function is multi-objective and reflects the quality of the generated mRNA according to three criteria:
  1. 1.
    GC-content: The fraction of G or C nucleotides in the sequence. High GC-content generally correlates with increased mRNA stability and translation efficiency (courel2019gc), in particular, optimizing GC-content has a similar effect to optimizing codon usage (zhang2023algorithm).
  2. 2.
    Minimum free energy (MFE): The (negative) folding free energy of the mRNA’s secondary structure. A lower MFE (more negative) indicates a more stable folded structure.
  3. 3.
    Codon adaptation index (CAI): A measure of how well the sequence conforms to species-specific codon usage preferences. Higher CAI typically leads to more efficient translation (sharp1987codon).
These components are linearly combined:
$R(s)=w_{1}\cdot\text{GC}(s)-w_{2}\cdot\text{MFE}(s)+w_{3}\cdot\text{CAI}(s),$
where $w\in\mathbb{R}^{3}$ are user-defined weights. This reflects the intuition from prior work that mRNA design must balance structural stability and codon optimality. However, these objectives are often in tension, making diversity and mode coverage essential for identifying high-quality trade-offs in the design space. Across the three GFlowNet variants, TB, SAGFN, and LGGFN, we generated diverse mRNA sequences encoding the same target protein. Despite the shared amino acid sequence, the models produced variants that varied significantly in their codon usage, GC content, and predicted RNA stability by the MFE metric. Notably, LGGFN consistently yielded sequences with high Codon Adaptation Index (CAI) and GC content, reflecting improved translation potential and structural stability. Furthermore, analysis of pairwise Levenshtein distances between generated sequences revealed substantial diversity, especially in LGGFN and SAGFN outputs. This diversity suggests that the models are exploring multiple high-reward regions of the vast mRNA sequence space.